2. Assume that the economy of Stockland produces four goods, rocks, socks, blocks, and clocks. Rocks and blocks are both used as weapons; they are good substitutes for each other. The quantities and prices for each of the goods in years one and two are given by the following table:

                             YEAR 1                                                           YEAR 2               

               good           quantity           price                 good           quantity           price                  

               rocks            200                  $1                 rocks                150                $4

               socks            200                  $2                 socks                220                $5

               blocks          200                  $3                 blocks             400                $2

               clocks           200                  $4                 clocks              300                $6

            a. What is nominal GDP in year 1?

            b. What is real GDP in year 1?

          c. What is nominal GDP in year 2?

            d. What is real GDP in year 2? (Use year 1 as the base year.)

            e. What is the percentage increase in nominal GDP? Real GDP?

            f. What is the GDP price deflator?

3. Using the data from problem 2), find the chain-type measures of GDP.

a. What is the rate of growth in real GDP using the chain-type technique (g)?  

b. What is the level of chain-type real GDP in year 2?

c. What is the chain-type GDP price deflator for year 2?

d. How do these compare with your answers from question 2)?
Answer #1

Nominal GDP = current year quantity * current year price

Real GDP = Current year quantity*base year price

a) NGDP in 1 = 200*1+200*2+200*3+200*4 = 200+400+600+800 = 2000

b) RGDP in 1 = same as NGDP = 2000

c) NGDP in 2 = 150*4+220*5+400*2+300*6 = 600+1100+800+1800 = 4300

d) RGDP in 2 = 150*1+220*2+400*3+300*4 = 150+440+1200+1200 = 2990

e) Percentage change in NGDP = 4300-2000/2000*100 = 115%

Percentage change in RGDP = 2990-2000/2000*100 = 49.5%

f) GDP Deflator = NGDP/RGDP*100

Year 1 = 2000/2000*100 =100

Year 2 = 4300/2990*100 = 143.81

As per Chegg guidelines, the first question in answered